OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line shown in the graph. Please show all of your work. (-6, -6) (6, 4)

OpenStudy (bahrom7893):

Slope = (y2-y1)/(x2-x1) = (4-(-6))/(6-(-6)) = 10/12 = 5/6

OpenStudy (bahrom7893):

equation of the line: y-y1 = M(x-x1) y-(-6) = (5/6)(x-(-6)) y+6 = (5/6)(x+6) y = (5/6)x + 5 - 6 y = (5/6)x - 1